# A Stochastic Approach to the Definition of the Path Integral Measure

### v2

@@ 1 Introduction
In the present work, we provide a possible approach to resolving the issues of convergence and dimensionality by restricting the relevant geometry to a neighborhood of the classical trajectory, which through fibration also provides an avenue for establishing a Gaussian measure for integration, yielding a controlled, well-defined functional integral over a stochastic process. Thus, the difficulties of the path integral stem not only from oscillatory factors but from the choice of path space itself, particularly the ill-defined “sum over all paths” which includes trajectories that are physically inaccessible to the system. Classical mechanics already suggests strong restrictions: for a fixed energy, the Euler–Lagrange equations determine a unique trajectory between two points on a configuration manifold. Even when the energy varies within a small uncertainty window, admissible paths remain confined to a bounded tubular neighborhood of the classical trajectory. Extrapolating this to the quantum domain, the finite spread of a wave packet and the uncertainty relation imply that only trajectories within a controlled deviation from the classical path contribute significantly. Thus, a physically meaningful reformulation of the path integral may be built on a restricted path space $C([0,1],M)$ that encompasses all physically relevant (accessible) trajectories.
@@
In the present paper, we attempt to develop a rigorous stochastic framework on such a restricted path space. By constructing a tubular neighborhood and expressing deviations as sections of the normal bundle, we obtain a restricted path space on which integration proceeds in a well-defined manner. The view of the tubular neighborhood arises naturally. If one considers the alternative paths taken by a particle while traveling from A to B and accounts for the boundedness in maximal allowed deviation from the classical trajectory, imposed by finiteness of energy, then we obtain a trajectory with a peak at a given point whose magnitude we may denote $\eta$. However, the peak may be displaced along the whole classical trajectory, and even rotated, yielding a cylinder with radius $\eta.$ Within this geometric setting, we define a stochastic path integral as a Gaussian expectation weighted by the physical action, and we show that it admits a semigroup structure analogous to the Feynman–Kac representation.
@@
Finally, for full physical and mathematical equivalence, we prove that the Stochastic Path integral holds under Wick transformation and, in fact, is equivalent to the Feynman-Kac Path Integral, which allows us to carry over the vast array of known theorems and properties of Euclidean path integrals onto our stochastic framework. In the current paper, we will only view the case when 1) a classical trajectory is given with fixed endpoints, 2) the particle surely starts and ends at the endpoints of the classical trajectory, 3) any trajectory not ending at the endpoint of the classical trajectory does not exist, 4) the particle is non-relativistic. These conditions are technicalities imposed for simplicity of formulation: nothing in the the framework forbids extensions to the domains where those conditions fail, however that necessitates a lot more work and refinement for mathematical consistency and rigor.
@@ 2.1 The Classical View
$\forall E\in\sigma$ where $\sigma$ is the non-degenerate part of the energy spectrum for a physical system with the Lagrangian $L(q,\dot{q}),\quad\exists\gamma_{E}\in\mathcal{P},\gamma_{E}:[0,1]\rightarrow M,\gamma_{E}(0)=A,\gamma(1)=B.$ with $\gamma_{E}$ unique.
@@
The value of energy $E$ is given by the inverse-map from phase space from $M$ of $H\in C^{\infty}$, while $\Phi_{E}:M\rightarrow\mathbb{R}$ satisfies $H(q,\nabla W_{E}(q))=E$. The trajectory $\gamma_{E}$ that corresponds to the path taken by the system given it having energy $E$ is then obtained by integrating $\dot{q}^{i}(t)=\frac{\partial H}{\partial p_{i}}(q,\nabla W_{E}(q)),q(t_{0})=A.$ Given that $H$ is convex in $p$, the HJE determines a unique element in $C^{1}$ and hence ascribes a unique $\gamma$ to all energies. Since $E$ is non-degenerate, there isn’t a configuration of the system with the same energy but different ”state”. For example, if we have two objects with the same energy but one has some non-zero angular momentum, that second object is excluded from our analysis. ∎
The uniqueness of trajectories for any given fixed energy value is a rather trivial result, given that the equation $\frac{\partial f}{\partial x}=\frac{\partial g}{\partial y}$ is well-posed. Now, let us assume the case where the energy isn’t strictly fixed. Rather, the energy varies under some small uncertainty values. In such cases, we may consider a Lie Algebra of transformations of some base trajectory $\gamma_{0}$ under parameters $\alpha$. We may evaluate this through symplectic geometry on $T^{*}M$ (Phase Space).
@@ 2.2 Hamilton’s Principle and Boundedness
The Hamilton Principle state that $\delta S=0$, wherefrom $S(q+\delta q,\dot{q}+\delta\dot{q})-S(q,\dot{q})=0$. Thus, for any trajectory inside $\mathfrak{B}$, we may hypothesize the existence of some $\eta>0$ such that $|S(\gamma^{\alpha})-S(\gamma_{0})|\leq\eta$ for any trajectory that is sufficiently close yet not equal to $\gamma_{0}$ and some variance in energy of the system, which lies within $E-\delta E,E+\delta E]$.
@@ 2.3 The Quantum-Mechanical View
Let $\hat{H}\psi=\sigma\psi$. We will be viewing a case of a 1-dimensional particle which may be very easily generalized to a multi-dimensional case. In the semiclassical view, the ansatz $\psi(x,t)=A(x,t)\exp(\frac{i}{\hbar}S(x,t)):-\frac{\partial S}{\partial t}=E(x,t)=\sigma(x,t)$ near the classical trajectory $\gamma_{0}.$ Let $T=t_{2}-t_{1}$ be the physical lifetime of $\psi$’s evolution from $A$ to $B$. Then, for the stationary case, $S_{\sigma}(x,t)=\Phi_{\sigma}(x)-\sigma t$, so along $\gamma$ with energy $E$, $S_{E}(\gamma)=S_{\sigma}(\gamma)+\sigma T-ET\implies\delta S(\gamma)=S_{\sigma+\delta E}(\gamma)-S_{\sigma}(\gamma)=-\delta ET$, which corresponds to the change in action due to change in energy from $\sigma$ to $\sigma+\delta E$. Given the finiteness of $\delta ET,$ we already get the hint that some bound of form of $\eta$ exists.
@@
It thus follows that $|\delta S(\gamma)|=|S_{\sigma+\delta E}(\gamma)-S_{\sigma}(\gamma)|=|\delta E|T\leq\eta\implies|\delta E|=\frac{\eta}{T}$. Thus, $|E-\sigma|\leq\frac{\eta}{T}$. By the uncertainty principle, we get $\frac{1}{2}\hbar\leq\eta$, which gives us that the diameter of the tubular neighborhood is $\sqrt{2\hbar}$
@@
3 On Relevant Trajectories
$f(E)=\frac{1}{(\delta E)^{2}-(\Delta E)^{2}}\implies f(H)=\frac{1}{(\delta E)^{2}-(H-E_{0})^{2}}$ (3.1)
$\Phi(\gamma,\gamma_{0})=\oint_{\gamma\circ\gamma_{0}}f(H)p_{i}dq^{i}$ (3.2)
For any $\gamma:|H(\gamma)-E_{0}|>\delta E,$ the contour approaches a pole of $f(H)$, wherefrom by the Residue Theorem $\Phi(\gamma,\gamma_{0})$ diverges. The class of admissible trajectories is therefore given by the homotopies of $\gamma_{0}$ for which $\Phi$ is finite.
@@ 4.1 Geometry of the $\eta$-Tube and Bundle structure
$P_{\eta}$ is a complete over $C^{\infty}(M,[0,1])$ .
Proof: $P_{\eta}:=\{\gamma=\hom\gamma_{0}:\rho(\gamma,\gamma_{0})(x)\leq\eta\}$, such that the metric $\rho$ is defined as an equivalence relation between elements of $P_{\eta}$ and $\gamma_{0}$ is treated as the 0-element of $P_{\eta}$, with the latter closed under additive group $G_{+}$. It may be defined through the metric of $M$ as $\rho(\gamma,\gamma_{0})=\sup g(\gamma,\gamma_{0})$. Then, $\exists\gamma_{k}:\rho(\gamma,\gamma_{0})\leq\rho(\gamma,\gamma_{k})+\rho(\gamma_{k},\gamma_{0})\leq\eta$. However, $\rho(\gamma_{k},\gamma_{0})\leq\eta\implies\gamma_{k}\in P_{\eta},\rho(\gamma_{k},\gamma)<\varepsilon:\eta-\varepsilon=\rho(\gamma_{k},\gamma_{0}).$ Since $\gamma=\hom\gamma_{0},\forall x\in[0,1],\gamma_{0}(x)\leq\gamma_{k}(x)\leq\gamma(x),\gamma_{k}$ is in the homotopy class of $\gamma$ and hence (by transitivity) in the homotopy class of $\gamma_{0}$, i.e $\gamma_{k}\in P_{\eta}$. Hence, $\exists(\gamma_{k},\gamma_{i})\in P_{\eta}:\rho(\gamma_{k},\gamma_{i})|P<\varepsilon\implies P_{\eta}$ is complete.
@@
$\Omega[A,B]$ is a $C^{\infty}$ Banach Manifold on $H^{1}$. In the tubular neighborhood of $\gamma_{0}$ $C$, its associated chart is $\Phi\in C^{\infty}$ with inverse $\ell.$
Proof: $\forall x\in M,\exists N_{\mathcal{T}}$ is the normal neighborhood for which $\exp_{x}:T_{x}M\rightarrow M$ is a $C^{\infty}$ diffeomorphism. Taking all the points $x\in M:x\cap\gamma_{0}\neq\emptyset$, we obtain a family of diffeomorphisms $\exp_{\gamma_{0}(t)}:B_{\eta}(0)\subset T_{\gamma_{0}(t)}M\rightarrow M$ that are fiberwise mapping, which yield a normal bundle to the baseline trajectory. Therefore, the Riemann exponent on the local normal chart is invertible and well-defined. Let it be labeled $\Phi$ (exponent), and $\ell$ for the logarithm. $\forall\gamma\in\Omega$, we may formulate a local chart, each with injectivity radius $r<<\eta$, modeled on $H^{1}$ and of $C^{\infty}$ kind. The associated Sobolev space, however, is defined over $T_{\gamma_{0}(t)}M$, which is equivalent to $\gamma^{*}TM$ for all $t\in[0,1]$. Transition maps from each tubular neighborhood are compositions of the Riemann functions (exp and log), and hence also $C^{\infty}$. Since the Sobolev space is Banach by definition, $\Omega[A,B]$ is a Banach manifod, given that the chart map and inverse are defined, smooth, act as local diffeomorphisms fiberwise and transition maps are also smooth.
@@ 4.2 Stochastic Dynamics in $\mathcal{T}_{\eta}$
With the geometry established and the formulation of the basis for the stochastic dynamics around all the homotopies of $\gamma_{0}$ through fibration, we may consider a stochastic process in each flux space for modeling the random perturbations of homotopies of the classical trajectory. Through the stochastic dynamics in each fiber, we may obtain all possible paths that the particle may take. Let $\mathcal{T}_{\eta}(\gamma)\subset M$ denote the $\eta$–tube around $\gamma$, realized (for $\eta>0$ sufficiently small) as the image under the exponential map of the open disk bundle in the normal bundle $N\gamma\subset\gamma^{*}TM$, with the tube being given by

### v3

@@ 1 Introduction
In the present work, we provide a possible approach to resolving the issues of convergence and dimensionality by restricting the relevant geometry to a neighborhood of the classical trajectory, which through fibration also provides an avenue for establishing a Gaussian measure for integration, yielding a controlled, well-defined functional integral over a stochastic process. Thus, the difficulties of the path integral stem not only from oscillatory factors but from the choice of path space itself, particularly the ill-defined “sum over all paths” which includes trajectories that are physically inaccessible to the system. Classical mechanics already suggests strong restrictions: for a fixed energy, the Euler–Lagrange equations determine a unique trajectory between two points on a configuration manifold under certain boundary conditions (see Lemma 1). Even when the energy varies within a small uncertainty window, admissible paths remain confined to a bounded tubular neighborhood of the classical trajectory. Extrapolating this to the quantum domain, the finite spread of a wave packet and the uncertainty relation imply that only trajectories within a controlled deviation from the classical path contribute significantly. Thus, a physically meaningful reformulation of the path integral may be built on a restricted path space $C([0,1],M)$ that encompasses all physically relevant (accessible) trajectories.
@@
In the present paper, we attempt to develop a rigorous mathematical framework on such a restricted path space by considering a quadratic Lagrangian for a non-degenerate system and reformulating the trajectories through a stochastic process on bundles defined on a tubular neighborhood of a classical trajectory. By constructing a tubular neighborhood and expressing deviations as sections of the normal bundle, we obtain a restricted path space on which integration proceeds in a well-defined manner. The view of the tubular neighborhood arises naturally when considering a free particle. If one considers the alternative paths taken by a particle while traveling from A to B and accounts for the boundedness in maximal allowed deviation from the classical trajectory, imposed by finiteness of energy, then we obtain a trajectory with a peak at a given point whose magnitude we may denote $\eta$. However, the peak may be displaced along the whole classical trajectory, and even rotated, yielding a cylinder with radius $\eta.$ Within this geometric setting, we define a stochastic path integral as a Gaussian expectation weighted by the physical action, and we show that it admits a semigroup structure analogous to the Feynman–Kac representation.
@@
Finally, for full physical and mathematical equivalence, we prove that the Stochastic Path integral holds under Wick transformation and, in fact, reproduces the results of the Feynman-Kac Theorem, which allows us to carry over the vast array of known results and properties of Euclidean path integrals onto our stochastic framework. In the current paper, we will only view the case when 1) a classical trajectory is given with fixed endpoints, finite time and finite energy uncertainty 2) the particle surely starts and ends at the endpoints of the classical trajectory, 3) any trajectory not ending at the endpoint of the classical trajectory does not exist, 4) the particle is non-relativistic. These conditions are technicalities imposed for simplicity of formulation: nothing in the framework forbids extensions to the domains where those conditions fail, however that necessitates a lot more work and refinement for mathematical consistency and rigor. We will be analyzing the case of the Gaussian free particle in this paper, yet the stochastic structure leveraged for omitting trajectories explicitly in integration will be formulated as useful for the general case, accounting for drift and more. The evolution of the system accounting for energy uncertainty, degenerate trajectories and infinite timescales will not be explored.
@@ 2.1 The Classical View
$\forall E\in\sigma$ where $\sigma$ is the non-degenerate part of the energy spectrum for a physical system with a quadratic Lagrangian $L(q,\dot{q}),\quad\exists\gamma_{E}\in\mathcal{P},\gamma_{E}:[0,1]\rightarrow M,\gamma_{E}(0)=A,\gamma(1)=B.$ with $\gamma_{E}$ unique given $[0,T]$ contains no conjugate points and $T<\infty$.
@@
The value of energy $E$ is given by the inverse-map from phase space from $M$ of $H\in C^{\infty}$, while $\Phi_{E}:M\rightarrow\mathbb{R}$ satisfies $H(q,\nabla W_{E}(q))=E$. The trajectory $\gamma_{E}$ that corresponds to the path taken by the system given it having energy $E$ is then obtained by integrating $\dot{q}^{i}(t)=\frac{\partial H}{\partial p_{i}}(q,\nabla W_{E}(q)),q(t_{0})=A.$ Given that $H$ is convex in $p$, the HJE determines a unique element in $C^{1}$ and hence ascribes a unique $\gamma$ to all energies. Since $E$ is non-degenerate, there isn’t a configuration of the system with the same energy but different ”state”. For example, if we have two objects with the same energy but one has some non-zero angular momentum, that second object is excluded from our analysis. Since there are no conjugate points contained in the interval, $\delta^{2}S$ is at all times non-degenerate, alleviating the issue of local locus-straying (minimal deviations induced by the Jacobi field). Hence, uniqueness is guaranteed. ∎
For the entirety of the paper, we will be assuming systems under the conditions of Lemma 1. The uniqueness of trajectories for any given fixed energy value is a rather trivial result when discussing the boundary conditions imposed in Lemma 1, yet it is valuable to gain perspective into the nature of classical dynamics for non-degenerate systems. Importantly, unless the boundary conditions are exactly those imposed for Lemma 1, the plurality of trajectories becomes possible. Problems such as the non-uniqueness of trajectories for $T\rightarrow\infty$ and degenerate energies will not be explored in this paper for that would necessitate a global analysis. Now, let us assume the case where the energy isn’t strictly fixed, varying under some small uncertainty values. The uncertainty in energy will also be considered as finite, as the analysis of systems with infinite uncertainties again collapses to the case of studying the global evolution for the system over the entire energy spectrum, which is outside of the scope of the present paper. In our case, we may consider a Lie Algebra of transformations of some base trajectory $\gamma_{0}$ under parameters $\alpha$. We may evaluate this through symplectic geometry on $T^{*}M$ (Phase Space).
@@ 2.2 Hamilton’s Principle and Boundedness
The Hamilton Principle states that $\delta S=0$, wherefrom $S(q+\delta q,\dot{q}+\delta\dot{q})-S(q,\dot{q})=0$. Thus, for any trajectory inside $\mathfrak{B}$, we may hypothesize the existence of some $C>0$ such that $|S(\gamma^{\alpha})-S(\gamma_{0})|\leq C$ for any trajectory that is sufficiently close yet not equal to $\gamma_{0}$ and some variance in energy of the system, which lies within $E-\delta E,E+\delta E]$.
@@ 2.3 The Quantum-Mechanical View
Let $\hat{H}\psi=\sigma\psi$. We will be viewing a case of a 1-dimensional particle which may be very easily generalized to a multi-dimensional case. In the semiclassical view, the ansatz $\psi(x,t)=A(x,t)\exp(\frac{i}{\hbar}S(x,t)):-\frac{\partial S}{\partial t}=E(x,t)=\sigma(x,t)$ near the classical trajectory $\gamma_{0}.$ Let $T=t_{2}-t_{1}$ be the physical lifetime of $\psi$’s evolution from $A$ to $B$. Then, for the stationary case, $S_{\sigma}(x,t)=\Phi_{\sigma}(x)-\sigma t$, so along $\gamma$ with energy $E$, $S_{E}(\gamma)=S_{\sigma}(\gamma)+\sigma T-ET\implies\delta S(\gamma)=S_{\sigma+\delta E}(\gamma)-S_{\sigma}(\gamma)=-\delta ET$, which corresponds to the change in action due to change in energy from $\sigma$ to $\sigma+\delta E$. Given the finiteness of $\delta ET,$ we already get the hint that some bound of form of $C$ exists.
@@
It thus follows that $|\delta S(\gamma)|=|S_{\sigma+\delta E}(\gamma)-S_{\sigma}(\gamma)|=|\delta E|T\leq C\implies|\delta E|=\frac{C}{T}$. Thus, $|E-\sigma|\leq\frac{C}{T}$. Heuristically, by the uncertainty principle, we get $\frac{1}{2}\hbar\leq C$. Here, like in semiclassical analysis, the diameter of our tube will be on the order of $\sqrt{2\hbar}$
@@ 3.1 Geometry of the $\eta$-Tube and Bundle structure
Let $(M,g)$ be a Riemannian manifold with a fixed continuous reference path $\gamma_{0}\in C([0,1],M)$ defined over it. For $\eta>0$ and a supremum metric $\rho(x,y)=\sup d_{g}(x,y)$, $(P_{\eta},\rho)$ is complete.
Proof: The metric space $(C([0,1],M),\rho)$ is complete whenever $(M,d_{g})$ is complete. Let us parametrize $\rho$ by some $t\in[0,1]$, with $\rho(x,y)=\sup_{t\in[0,1]}d_{g}(x(t),y(t))$. If $\gamma_{k}$ is $\rho$–Cauchy, then $\forall t\in[0,1]$, the sequence $\gamma_{k}(t)$ is $d_{g}$–Cauchy and hence converges in $M$; the pointwise limit $\gamma$ is continuous and $\gamma_{k}\to\gamma$ uniformly. Now, let $\gamma_{k}\subset P_{\eta}$ be $\rho$–Cauchy. By completeness of $C([0,1],M)$, $\gamma_{k}\to\gamma$ in $\rho$ for some $\gamma\in C([0,1],M)$. Since the map $\gamma\rightarrow\rho(\gamma,\gamma_{0})$ is $\rho$–continuous, $\rho(\gamma,\gamma_{0})=\lim_{k\to\infty}\rho(\gamma_{k},\gamma_{0})\leq\eta,$ so $\gamma$ satisfies the tube constraint. By fixing $k$ and choosing some $r_{k}>0:r_{k}<\text{inj}_{M}(\gamma_{k}(t))$ for all $t\in[0,1]$. For all sufficiently large $j$ we have $\rho(\gamma_{j},\gamma_{k})<r_{k}$, hence for each $t$ there is a unique minimizing geodesic from $\gamma_{k}(t)$ to $\gamma_{j}(t)$ depending continuously on $t$ on M. This defines a homotopy $H_{k,j}:[0,1]\times[0,1]\to M$ under fixed endpoints from $\gamma_{k}$ to $\gamma_{j}$. Passing to the limit $j\to\infty$ yields a homotopy $H_{k}$ from $\gamma_{k}$ to $\gamma$. Since $\gamma_{k}\sim\gamma_{0}$ relative to endpoints by transitivity, and hence $\gamma\in P_{\eta}$. Therefore, every $\rho$–Cauchy sequence in $P_{\eta}$ converges in $(P_{\eta},\rho)$, i.e. $P_{\eta}$ is complete.
@@
On the manifold M, we may define a Sobolev space pinned by A,B as $\Omega[A,B]:=\{\gamma\in H^{1}([0,1],M):\ \gamma(0)=A,\ \gamma(1)=B\}.$ $\Omega[A,B]$ carries a natural structure of a $C^{\infty}$ Banach manifold modeled on $H^{1}_{0}([0,1],\mathbb{R}^{n}):$ $n=\dim M$. If $\gamma_{0}\in\Omega[A,B]$ and $\eta>0$ is such that $\exp_{\gamma_{0}(t)}$ is a diffeomorphism on $B_{\eta}(0)\subset T_{\gamma_{0}(t)}M$ for all $t$, then the tubular chart $\Phi:U_{\eta}\subset H^{1}_{0}([0,1],\gamma_{0}^{*}TM)\to\Omega[A,B],\qquad\Phi(v)(t):=\exp_{\gamma_{0}(t)}(v(t)),$ is $C^{\infty}$ with inverse $\ell:\Phi(U_{\eta})\to H^{1}_{0}([0,1],\gamma_{0}^{*}TM),\qquad\ell(\gamma)(t):=\exp_{\gamma_{0}(t)}^{-1}(\gamma(t)).$
Proof: We will again define everything relative to some $\gamma_{0}\in\Omega[A,B]$. By continuity of $t\rightarrow\gamma_{0}(t)$ and positivity of the injectivity radius along compact sets (which our intervals are, for example), $\exists\eta>0$ such that for every $t\in[0,1]$ the exponential map $\exp_{\gamma_{0}(t)}:B_{\eta}(0)\subset T_{\gamma_{0}(t)}M\to M$ is a $C^{\infty}$ diffeomorphism. Let $\gamma_{0}^{*}TM\to[0,1]$ be the pullback bundle and set $U_{\eta}:=\{v\in H^{1}_{0}([0,1],\gamma_{0}^{*}TM):\ \|v\|_{L^{\infty}}<\eta\}.$ Since $H^{1}([0,1],\mathbb{R}^{n})\hookrightarrow C^{0}([0,1],\mathbb{R}^{n})$, the $L^{\infty}$-constraint makes $U_{\eta}$ open in $H^{1}_{0}([0,1],\gamma_{0}^{*}TM)$. Then, let $\Phi$ be defined such that $\Phi(v)(t)=\exp_{\gamma_{0}(t)}(v(t))$. Then $\Phi(v)\in H^{1}([0,1],M)$ and $\Phi(v)(0)=A$, $\Phi(v)(1)=B$ because $v(0)=v(1)=0$, and hence $\Phi:U_{\eta}\to\Omega[A,B]$ is well-defined. In any local trivialization of $\gamma_{0}^{*}TM$ over $[0,1]$, the map $(t,w)\rightarrow\exp_{\gamma_{0}(t)}(w)$ is $C^{\infty}$ in $(t,w)$ on $0,1]\times B_{\eta}(0)$, which by Nemytskii theorems in 1D Sobolev spaces (composition theorems) [13], the induced map $v\mapsto\Phi(v)$ is $C^{\infty}$ as a map between Banach spaces $H^{1}_{0}\to H^{1}$. Since $\exp_{\gamma_{0}(t)}$ is a diffeomorphism on $B_{\eta}(0)$, the pointwise inverse $\exp_{\gamma_{0}(t)}^{-1}$ yields $\ell(\Phi(v))=v$ and $\Phi(\ell(\gamma))=\gamma$ on $\Phi(U_{\eta})$, and the same composition theorem gives that $\ell$ is $C^{\infty}$. For a general $\gamma\in\Omega[A,B]$, we may repeat the construction with $\gamma$ in place of $\gamma_{0}$ to obtain a family of charts modeled on $H^{1}_{0}([0,1],\gamma^{*}TM)$. On overlaps of the charts, transition maps are of the form $v\rightarrow\exp_{\gamma(t)}^{-1}\!\bigl(\exp_{\gamma^{\prime}(t)}(v(t))\bigr)$ and hence are $C^{\infty}$ by the same Sobolev composition argument. Therefore, $\Omega[A,B]$ is a $C^{\infty}$ Banach manifold, and in the tubular neighborhood of $\gamma_{0}$ the chart is $\Phi$ with inverse $\ell$. ∎
@@ 3.2 Stochastic Dynamics in $\mathcal{T}_{\eta}$
In the present section and onwards, we will concern ourself with formulating the stochastic dynamics through which the various trajectories the particle’s evolution occurs over may be defined without an explicit summation over them. We will formulate it in a general way that for a free particle may be seen as excessive. This section may be skipped entirely, as section 4 and 6 concern themselves purely with that which may be used for the free particle case. Naturally, for systems of a different sort–harmonic oscillators, particles in potential fields, etc–cannot be defined in a trivial tubular way that a free particle can, as the path space corresponding to those systems requires a different sort of compactification. Nevertheless, what applies for the stochastic dynamics determining trajectories here may be carried over to the general case, although possibly with some modifications.
We may consider now a stochastic process in each flux space for modeling the random perturbations of homotopies of the classical trajectory. Through the stochastic dynamics in each fiber, we may obtain all possible paths that the particle may take. Let $\mathcal{T}_{\eta}(\gamma)\subset M$ denote the $\eta$–tube around $\gamma$, realized (for $\eta>0$ sufficiently small) as the image under the exponential map of the open disk bundle in the normal bundle $N\gamma\subset\gamma^{*}TM$, with the tube being given by